# One-loop contributions to $WWZ$ from new physics in the neutrino sector

### v1

@@ II Theoretical aspects
Throughout this Section, we discuss the neutrino model given in Ref. Pilaftsis , which is the framework for our phenomenological calculation. General aspects regarding the AC which we wish to calculate are addressed as well.
II.1 Majorana neutrinos in physics beyond the SM
The neutrino sector of the SM is built upon the assumption that neutrinos are properly described by massless and electrically-neutral Dirac spinor fields. So far, no electromagnetic interaction of a neutrino has ever been measured, thus no evidence in contraposition to the electromagnetic innocuity of these particles exists. Note, however, that works assuming millicharged neutrinos are available DHR ; StTo . Another NP model, the so-called “minimally extended SM” GiKi , which is given by the introduction of three right-handed neutrino fields, together with a set of neutrino Yukawa-sector terms, produces a contribution to the neutrino magnetic moment [^2] For a review on neutrino electromagnetic properties, including an estimation of the magnetic moment of the neutrino, see BGS and references therein. of order $10^{-19}\,\mu_{\rm B}$, with $\mu_{\rm B}$ the Bohr magneton, which lies 77~{}77 orders of magnitude beyond current experimental sensitivity XENONnuMM . Throughout the present paper, neutrinos are assumed to be electrically neutral. Neutrino oscillations, according to which the probability of observing a neutrino flavor different from the one originally created at some source is nonzero, were first measured by the Super-Kamiokande Collaboration Kamiokande , being corroborated, just four years later, at the Sudbury Neutrino Observatory SNO . About the massiveness of neutrinos, the phenomenon of neutrino oscillations imposes the restriction that at least two of them have nonzero masses Pontecorvo . Neutrino mixing, also required by neutrino oscillations, is characterized by the Pontecorvo-Maki-Nakagawa-Sakata (PMNS) neutrino-mixing matrix MNSmatrix ; Pontecorvomatrix , here denoted by $U_{\nu}$, which is conventionally parametrized by three mixing angles, denoted by $\theta_{12}$, $\theta_{23}$, and $\theta_{13}$, and up to three complex phases. The first measurements of the last mixing angle, $\theta_{13}$, were reported, almost simultaneously, by the Daya-Bay Collaboration DayaBay and the Reno Collaboration Reno , after which neutrino oscillations were fully accepted as a genuine physical phenomenon. One of the complex phases of $U_{\nu}$, known as the “Dirac phase”, has been addressed by the T2K Collaboration T2K1 ; T2K2 . Even though nowadays massiveness of neutrinos is widely accepted, such masses have never been measured, while experimental limits exist. Notably, the KATRIN Collaboration recently reported the upper bound $m_{\nu_{j}}<0.8\,{\rm eV}$, on neutrino masses KATRIN . Besides the goal of measuring neutrino masses, the relevant issue of the mechanism behind neutrino massiveness remains an open question. Though the usual SM way of generating fermion masses offers an answer, in which the size of the masses of neutrinos are decisively determined by the values of Yukawa constants, the unique smallness of such masses, in comparison with the mass spectrum of the remaining SM fermions, is quite noticeable. This has motivated a pursuit for alternative mass-generating mechanisms, not relying on ad hoc values of Yukawa constants, but rather being connected to some sort of NP. The construction of neutrino sectors, aimed at the determination of the origin of neutrino mass, depends on whether the neutrinos are assumed to be described by either Dirac Dirac or Majorana Majorana fields. A main difference among these descriptions is that in the case of Majorana fields the relation $\psi^{\rm c}=\psi$, with $\psi^{\rm c}=C\overline{\psi}^{\rm T}$ the charge-conjugated field of $\psi$ and $C$ the charge-conjugation matrix, holds. This equation, known as the Majorana condition, can be fulfilled only if $\psi$ is electrically neutral. Since neutrinos satisfy this requirement, the question of whether their description is correctly achieved by Dirac or Majorana fermions is pertinent and the determination of its answer currently incarnates a main objective in neutrino physics. The most representative physical process to look for Majorana neutrinos is the neutrinoless double beta decay, whose observation would be evidence of the Majorana nature of the neutrinos. A plethora of experimental collaborations have been working for several years, looking for a measurement of this physical process CUPIDMOndbd ; CUOREndbd ; GERDAndbd ; Majoranandbd ; EXO200ndbd ; KamLANDZenndbd , which, however, has never been observed nohayndbd , thus indicating that, even if neutrinos turn out to be associated to Majorana fields, the neutrinoless double beta decay is quite rare in nature. Alternative means to determine the Dirac or Majorana nature of neutrinos exist, such as the one given in Ref. MRS . The model of Ref. Pilaftsis , used for the present investigation, assumes the neutrinos to be Majorana-like.
Effective field theories are model-independent descriptions, built upon low-energy dynamic variables and symmetries, which provide generic parametrizations of effects associated to unknown physics characterized by dynamic variables and symmetries governing nature at some high-energy scale $\Lambda$ DGMP . When the low-energy formulation is renormalizable, as in the case of the SM, each of the corresponding effective-Lagrangian terms, with mass units $n>4$, involves a suppression factor $\frac{\alpha_{\rm eff}}{\Lambda^{n-4}}$, where the $(n-4)$-th power of the high-energy scale $\Lambda$ corrects the units of the Lagrangian term to render them $4$, whereas $\alpha_{\rm eff}$ is a dimensionless parameter determined by the high-energy theory itself. So, even though effective theories comprise an infinite number of Lagrangian terms, thus entailing an infinite number of parameters, such terms are hierarchical, which allows for the consideration of some finite subset of the whole effective theory. While the derivation of an effective Lagrangian can be achieved by integrating out heavy degrees of freedom from some given model, their usage is usually way more pragmatic. If no specific high-energy description is considered, the low-energy dynamic variables and symmetries can be used to construct a subset of the infinite set of nonrenormalizable effective-Lagrangian terms. In particular, the SM fields and its symmetries have been used to construct the whole set of effective-Lagrangian terms with units $\big{(}{\rm mass}\big{)}^{6}$ LLR ; BuWy . The low-energy symmetry criteria for the construction of effective-Lagrangian terms are often relaxed in order to look for NP beyond the SM. This is the case of the Weinberg operator Weinbergoperator , a mass-dimension-5 effective-Lagrangian term which, by contrast with the SM, violates lepton-number symmetry and which is given by ${\cal L}_{\rm W}=-\frac{\kappa_{\alpha\beta}}{\lambda}\big{(}\overline{L^{\rm c}_{\alpha,L}\tilde{\phi}^{*}}\big{)}\big{(}\tilde{\phi}^{\dagger}L_{\beta,L}\big{)}$, where $L_{\alpha,L}$ is the $\alpha$-th left-handed ${\rm SU}(2)_{L}$ lepton doublet, $\phi$ is the Higgs doublet, and the definition $\tilde{\phi}=i\sigma^{2}\phi^{*}$ holds, with $\sigma^{2}$ the imaginary Pauli matrix. This effective theory requires neutrinos to be characterized by Majorana fields. The Weinberg operator yields, via the Brout-Englert-Higgs mechanism EnBr ; Higgs , Majorana neutrino-mass terms, with neutrino masses given as $m_{\nu_{j}}\sim\frac{v^{2}}{\Lambda}$, where $v=246\,{\rm GeV}$ is the SM Higgs field vacuum expectation value. Though, as any effective theory, the Weinberg operator does not fully specifies what the genuine high-energy description is, it provides restrictions on the allowed ultraviolet completions Ma . The current KATRIN bound on neutrino masses KATRIN , in the sub-eV scale, yields an estimation $\Lambda\sim 10^{13}\,{\rm GeV}$, for the high-energy scale.
@@ II.1 Majorana neutrinos in physics beyond the SM
Consider some NP theory, beyond the Standard Model (BSM), characterized by the Lagrangian density ${\cal L}_{\rm BSM}$. We assume that this formulation goes through a couple of stages of spontaneous symmetry breaking, the first of which takes place at some high energy scale, $\Lambda$, then resulting in a theory invariant with respect to the SM gauge symmetry group, ${\rm SU}(3)_{C}\otimes{\rm SU}(2)_{L}\otimes{\rm U}(1)_{Y}$. Later, at the energy scale $v$, the Brout-Englert-Higgs mechanism occurs, after which the electroweak group ${\rm SU}(2)_{L}\otimes{\rm U}(1)_{Y}$ is broken down into the electromagnetic gauge group, ${\rm U}(1)_{e}$. We assume that, at this point, the ${\cal L}_{\rm BSM}$ Lagrangian reads
@@
In the last equation, the presence of the four Lagrangian terms ${\cal L}^{\nu}_{\rm mass}$, ${\cal L}_{W\nu l}$, ${\cal L}_{Z\nu\nu}$, and ${\cal L}_{h\nu\nu}$ is explicitly indicated, whereas further Lagrangian terms, most of them characteristic of the NP specifically considered, are implied by the ellipsis.
The Lagrangian term ${\cal L}^{\nu}_{\rm mass}$ includes all the terms quadratic in neutrino fields. We assume this Lagrangian term to comprise six chiral neutrino fields, three of them left handed, here denoted by $\nu^{0}_{j,L}$, and the remaining three, represented as $\nu^{0}_{k,R}$, being right handed. Now we assume the Lagrangian term ${\cal L}^{\nu}_{\rm mass}$ to be given by
@@
$\hskip 34.14322pt+\frac{1}{2}\overline{\nu^{0{\rm c}}_{j,R}}(m_{\rm M})_{jk}\nu^{0}_{k,R}\Big{)}+{\rm H.c.}$ (5)
Here, $\nu^{0{\rm c}}_{j,L}$ and $\nu^{0{\rm c}}_{k,R}$ are the charge-conjugated fields of $\nu^{0}_{j,L}$ and $\nu^{0}_{k,R}$, respectively. Recall that implementation of charge conjugation to any fermion field with definite chirality yields a field with opposite chirality, so $\nu^{0{\rm c}}_{j,L}$ is right handed and $\nu^{0{\rm c}}_{k,R}$ is left handed. The terms in ${\cal L}^{\nu}_{\rm mass}$ involving the $3\times 3$ symmetric matrix $m_{\rm M}$, which mixes neutrino fields $\nu^{0{\rm c}}_{j,R}$ and $\nu^{0}_{k,R}$, resemble, upon some diagonalization procedure, a set of Majorana mass terms. We assume $m_{\rm M}$ to be generated by the first stage of symmetry breaking, at $\Lambda$, thus meaning that a connection among such a scale and this matrix exist. Furthermore, Dirac-like mixings of neutrino fields $\nu^{0}_{j,L}$ and $\nu^{0}_{k,R}$, driven by the $3\times 3$ matrix $m_{\rm D}$, also take place in ${\cal L}^{\nu}_{\rm mass}$. We think of the matrix $m_{\rm D}$ as linked to the vacuum expectation value $v$, through electroweak symmetry breaking. Moving on towards the definition of the basis of neutrino mass eigenspinors, the neutrino-mass Lagrangian is rearranged as
${\cal L}^{\nu}_{\rm mass}=-\frac{1}{2}\big{(}\overline{f_{L}}\,\,\,\,\,\overline{F_{L}}\big{)}{\cal M}\left(\begin{array}[]{c}f_{R}\\ F_{R}\end{array}\right)+{\rm H.c.},$ (6)
for which the $3\times 1$ matrices
$f_{L}=\left(\begin{array}[]{c}\nu^{0}_{1,L}\\ \nu^{0}_{2,L}\\ \nu^{0}_{3,L}\end{array}\right),\hskip 14.22636pt$ $F_{L}=\left(\begin{array}[]{c}\nu^{0{\rm c}}_{1,R}\\ \nu^{0{\rm c}}_{2,R}\\ \nu^{0{\rm c}}_{3,R}\end{array}\right),$ (13)
$f_{R}=\left(\begin{array}[]{c}\nu^{0{\rm c}}_{1,L}\\ \nu^{0{\rm c}}_{2,L}\\ \nu^{0{\rm c}}_{3,L}\end{array}\right),\hskip 14.22636pt$ $F_{L}=\left(\begin{array}[]{c}\nu^{0}_{1,R}\\ \nu^{0}_{2,R}\\ \nu^{0}_{3,R}\end{array}\right),$ (20)
have been defined, whereas the $6\times 6$ matrix ${\cal M}$ is given, in terms of $3\times 3$ matrix blocks, by
${\cal M}=\left(\begin{array}[]{cc}0&m_{\rm D}\\ m_{\rm D}^{\rm T}&m_{\rm M}\end{array}\right).$ (21)
Since $m_{\rm M}$ is symmetric, the property ${\cal M}^{\rm T}={\cal M}$ holds, which guarantees the existence of a diagonalization unitary matrix Takagi ,
${\cal U}=\left(\begin{array}[]{cc}{\cal U}_{11}&{\cal U}_{12}\\ {\cal U}_{21}&{\cal U}_{22}\end{array}\right),$ (22)
@@
${\cal U}^{\rm T}{\cal M}\,\,{\cal U}=\left(\begin{array}[]{cc}M_{\nu}&0\\ 0&M_{N}\end{array}\right),$ (23)
where $M_{\nu}$ and $M_{N}$ are $3\times 3$ diagonal real matrices, whose eigenvalues are strictly positive. The diagonalization displayed in Eq. (23) defines the change of basis given by
$\left(\begin{array}[]{c}\nu_{L}\\ N_{L}\end{array}\right)={\cal U}^{\rm T}\left(\begin{array}[]{c}f_{L}\\ F_{L}\end{array}\right),\hskip 11.38092pt\left(\begin{array}[]{c}\nu_{R}\\ N_{R}\end{array}\right)={\cal U}^{\dagger}\left(\begin{array}[]{c}f_{R}\\ F_{R}\end{array}\right),$ (24)
where $\nu_{1}$, $\nu_{2}$, and $\nu_{3}$ are the mass-eigenspinor light-neutrino fields, whereas $N_{1}$, $N_{2}$, and $N_{3}$ stand for the mass-eigenspinor heavy-neutrino fields. The entries in the diagonal of the matrix $M_{\nu}$, which we denote as $m_{\nu_{j}}$ for $j=1,2,3$, are identified as the masses of the light neutrinos $\nu_{j}$. On the other hand, $M_{N}$, with diagonal entries $m_{N_{j}}$ for $j=1,2,3$, characterizes the masses of the three heavy neutrinos $N_{j}$. Let us emphasize that the mass eigenspinor fields $\nu_{j}$ and $N_{j}$ satisfy the Majorana condition, that is, $\nu_{j}^{\rm c}=\nu_{j}$ and $N_{j}^{\rm c}=N_{j}$, so they are Majorana neutrino fields. In this basis, the neutrino-mass Lagrangian is finally expressed as
${\cal L}^{\nu}_{\rm mass}=\sum_{j=1}^{3}\Big{(}-\frac{1}{2}m_{\nu_{j}}\overline{\nu_{j}}\nu_{j}-\frac{1}{2}m_{N_{j}}\overline{N_{j}}N_{j}\Big{)}.$ (25)
As established above, the matrix $m_{\rm M}$ is assumed to emerge from the high-energy spontaneous symmetry breaking, at the scale $\Lambda$, that leads to the SM gauge group, whereas the assumption is made that the matrix $m_{\rm D}$ is generated by electroweak symmetry breaking, thus being connected to the electroweak scale $v$. With this in mind, let us assume that $m_{\rm M}\sim\Lambda$ and $m_{\rm D}\sim v$, and then, for a moment, consider a scenario in which $\Lambda\gg v$ holds. Under such circumstances, the light-neutrino mass matrix is $M_{\nu}\sim\frac{v^{2}}{\Lambda}$, which matches the neutrino masses generated by the Weinberg operator, and the corresponding mass matrix for heavy neutrinos turns out to be $M_{N}\sim\Lambda$. This is the seesaw neutrino-mass profile, in the context of the so-called type-1 seesaw mechanism. Originally meant to explain parity violation as a result of spontaneous symmetry breaking, left-right symmetric models engendered the seesaw mechanism MoSe1 ; MoSe2 ; PaSa . In this manner, the seesaw mechanism is able to explain tiny light-neutrino masses as an interplay with the set of masses of heavy partners, mediated by the high-energy scale $\Lambda$, thus avoiding the need for unnaturally small Yukawa constants. Nonetheless, this strong link among the masses of light and heavy neutrinos, created by the seesaw mechanism, implies that heavy-neutrino masses are huge, for, as previously discussed, $\Lambda\sim 10^{13}\,{\rm GeV}$, by virtue of the stringent upper bounds on the masses of light neutrinos. NP associated to an energy scale so large is expected, on general grounds, to generate quite suppressed effects at low energies, well beyond current experimental sensitivity.
@@
Aiming at the possibility of probing beyond-SM NP presumably responsible of generating the tiny light-neutrino masses, variants of the seesaw mechanism have been formulated. For instance, the inverse seesaw MoVa ; GoVa ; DeVa introduces, besides three heavy Majorana neutrinos, a further set of neutral fermions, also of Majorana type. Then, assumptions on quadratic neutral-fermion couplings yield a mass matrix which, written in terms of matrix blocks, resembles, before diagonalization, the one corresponding to the type-one seesaw mechanism. Further assumptions on the couplings define block matrices, within the mass matrix, constituted by tiny parameters, which are able to attenuate the pressure on the scale $\Lambda$ and, consequently, induce effects hopefully attainable by experimental facilities. The phenomenological calculations performed in the present investigation take as theoretical framework another seesaw variant, given in Ref. Pilaftsis , which we discuss. In that work, the conditions $({\cal M}{\cal U})_{jk}=0$, for $j=1,2,3,4,5,6$, are introduced as sufficient and necessary for the $k$-th light neutrino to be massless at tree level. In this manner, the mass terms for light neutrinos are eliminated from ${\cal L}^{\nu}_{\rm mass}$, though keep in mind that such assumptions do not mess with heavy-neutrino masses, which remain nonzero at tree level. By this mean, the tie originally defined by the seesaw mechanism to link the masses of light neutrinos to the masses of their heavy counterparts has been broken. Light neutrinos, known to be massive in accordance with the occurrence of neutrino oscillations, do get massive, but their masses follow from radiative corrections. Ref. Pilaftsis provides a calculation of one-loop masses of light neutrinos, from which a remarkable statement is made: a new link between light- and heavy-neutrino masses is established, according to which the former are tiny enough as long as the heavy-neutrino mass spectrum is quasi-degenerate.
@@
Here, $\alpha=e,\mu,\tau$, that is, this greek index labels the flavors of the SM leptons. The $3\times 3$ matrix $V^{\ell}$ is analogous to the Kobayahsi-Maskawa quark-mixing matrix KoMa , but for the lepton sector, though note that $V^{\ell}$ is not necessarily unitary. The set of quantities ${\cal B}_{\alpha\nu_{j}}$ and ${\cal B}_{\alpha N_{j}}$, just defined in Eqs. (26) and (27), respectively define the $3\times 3$ matrices ${\cal B}_{\nu}$ and ${\cal B}_{N}$, which can be accommodated as matrix blocks into the $3\times 6$ matrix
${\cal B}=\big{(}{\cal B}_{\nu}\,\,\,{\cal B}_{N}\big{)}.$ (28)
@@
${\cal B}_{\alpha j}=\left\{\begin{array}[]{lcl}{\cal B}_{\alpha\nu_{k}}\textrm{, if }j=1,2,3,\\ {\cal B}_{\alpha N_{k}}\textrm{, if }j=4,5,6,\end{array}\right.$ (29)
@@
${\cal C}=\left(\begin{array}[]{cc}{\cal C}_{\nu\nu}&{\cal C}_{\nu N}\\ {\cal C}_{N\nu}&{\cal C}_{NN}\end{array}\right).$ (32)
@@
${\cal C}_{jk}=\left\{\begin{array}[]{ll}{\cal C}_{\nu_{i}\nu_{l}},&\textrm{ if}j=1,2,3\textrm{ and }k=1,2,3,\\ {\cal C}_{\nu_{i}N_{l}},&\textrm{ if }j=1,2,3\textrm{ and }k=4,5,6,\\ {\cal C}_{N_{i}\nu_{l}},&\textrm{ if }j=4,5,6\textrm{ and }k=1,2,3,\\ {\cal C}_{N_{i}N_{l}},&\textrm{ if }j=4,5,6\textrm{ and }k=4,5,6,\end{array}\right.$ (33)
@@
Returning to the Lagrangian ${\cal L}_{\rm BSM}$, as expressed in Eq. (4), it explicitly displays, besides the neutrino-mass Lagrangian term ${\cal L}^{\nu}_{\rm mass}$, three further terms, namely, ${\cal L}_{W\nu l}$, which involves charged currents, ${\cal L}_{Z\nu\nu}$, consisting of neutral currents featuring the SM $Z$ boson, and ${\cal L}_{h\nu\nu}$, which comprehends the couplings of the Higgs boson with neutrinos. The couplings constituting ${\cal L}_{W\nu l}$ and ${\cal L}_{Z\nu\nu}$ were recently utilized to explore the contributions, at one loop, from Majorana neutrinos, both light and heavy ones, to the triple gauge vertices $WW\gamma$ MMNS and $ZZZ^{*}$ NoSa , in the context of the seesaw variant of Ref. Pilaftsis . The explicit expressions of the Lagrangian terms read
@@
${\cal U}=\left(\begin{array}[]{cc}\big{(}{\bf 1}_{3}+\xi^{*}\xi^{\rm T}\big{)}^{-\frac{1}{2}}&\xi^{*}\big{(}{\bf 1}_{3}+\xi^{\rm T}\xi^{*}\big{)}^{-\frac{1}{2}}\\ -\xi^{\rm T}\big{(}{\bf 1}_{3}+\xi^{*}\xi^{\rm T}\big{)}^{-\frac{1}{2}}&\big{(}{\bf 1}_{3}+\xi^{\rm T}\xi^{*}\big{)}^{-\frac{1}{2}}\end{array}\right).$ (42)
@@
${\cal C}=\left(\begin{array}[]{cc}\big{(}{\bf 1}_{3}+\xi\xi^{\dagger}\big{)}^{-1}&\big{(}{\bf 1}_{3}+\xi\xi^{\dagger}\big{)}^{-1}\xi\\ \xi^{\dagger}\big{(}{\bf 1}_{3}+\xi\xi^{\dagger}\big{)}^{-1}&\xi^{\dagger}\big{(}{\bf 1}_{3}+\xi\xi^{\dagger}\big{)}^{-1}\xi\end{array}\right).$ (44)

### v2

@@ Appendix B The neutrino model
Throughout the present Appendix, we discuss the neutrino-mass model given in Ref. Pilaftsis , which is the framework for our phenomenological calculation.
@@
Assume that a theory beyond the Standard Model (BSM) is subjected to a couple of stages of spontaneous symmetry breaking, first at some high-energy scale $\Lambda$ and then at $v=246\,{\rm GeV}$, after which the corresponding Lagrangian, gauge invariant under ${\rm U}(1)_{e}$, is written as
@@
The neutrino-mass Lagrangian, ${\cal L}^{\nu}_{\rm mass}$, given by
@@
$\hskip 34.14322pt+\frac{1}{2}\overline{\nu^{0{\rm c}}_{j,R}}(m_{\rm M})_{jk}\nu^{0}_{k,R}\Big{)}+{\rm H.c.},$ (37)
comprises three left-handed neutrino fields $\nu^{0}_{j,L}$ and three right-handed neutrino fields $\nu^{0}_{j,R}$, together with the charge-conjugated spinor fields $\nu^{0{\rm c}}_{j,R}$. The $3\times 3$ lepton-number-violating matrix $m_{\rm M}$ is assumed to originate form the symmetry breaking at $\Lambda$, whereas the Dirac-like matrix $m_{\rm D}$, $3\times 3$-sized, is assumed to come from electroweak symmetry breaking. Then, ${\cal L}_{\rm mass}^{\nu}$ is rearranged as
${\cal L}^{\nu}_{\rm mass}=-\frac{1}{2}\big{(}\overline{f_{L}}\,\,\,\,\,\overline{F_{L}}\big{)}{\cal M}\left(\begin{array}[]{c}f_{R}\vspace{0.2cm}\\ F_{R}\end{array}\right)+{\rm H.c.},$ (38)
where $(f_{L})_{j}=\nu^{0}_{j,L}$, $(F_{L})_{j}=\nu_{j,R}^{0{\rm c}}$, $(f_{R})_{j}=\nu_{j,L}^{0{\rm c}}$, and $(F_{R})_{j}=\nu^{0}_{j,R}$. Moreover,
${\cal M}=\left(\begin{array}[]{cc}0&m_{\rm D}\vspace{0.2cm}\\ m_{\rm D}^{\rm T}&m_{\rm M}\end{array}\right).$ (39)
Since $m_{\rm M}$ is symmetric, the property ${\cal M}^{\rm T}={\cal M}$ holds, which guarantees the existence of a diagonalization unitary matrix Takagi
${\cal U}=\left(\begin{array}[]{cc}{\cal U}_{11}&{\cal U}_{12}\vspace{0.2cm}\\ {\cal U}_{21}&{\cal U}_{22}\end{array}\right),$ (40)
@@
${\cal U}^{\rm T}{\cal M}\,\,{\cal U}=\left(\begin{array}[]{cc}M_{\nu}&0\vspace{0.2cm}\\ 0&M_{N}\end{array}\right),$ (41)
@@
where $M_{\nu}$ and $M_{N}$ are $3\times 3$ diagonal real matrices, whose eigenvalues are, respectively, the three-level light- and heavy-neutrino masses. The diagonalization procedure induces a change of basis, which defines three massless neutrino fields $\nu_{j}$ and three heavy-neutrino fields $N_{j}$, all of them fulfilling the Majorana condition. In Ref. Pilaftsis , the conditions $({\cal M}{\cal U})_{jk}=0$, for $j=1,2,3,4,5,6$, are introduced as sufficient and necessary for the $k$-th light neutrino to be massless at tree level. In this manner, the mass terms for light neutrinos are eliminated from ${\cal L}^{\nu}_{\rm mass}$, though keep in mind that such assumptions do not mess with heavy-neutrino masses, which remain nonzero at tree level. By this mean, the tie originally defined by the seesaw mechanism to link the masses of light neutrinos to the masses of their heavy counterparts has been broken. Light neutrinos, known to be massive in accordance with the occurrence of neutrino oscillations, do get massive, but their masses follow from radiative corrections. Ref. Pilaftsis provides a calculation of one-loop masses of light neutrinos, from which a remarkable statement is made: a new link between light- and heavy-neutrino masses is established, according to which the former are tiny enough as long as the heavy-neutrino mass spectrum is quasi-degenerate.
@@
Here, $\alpha=e,\mu,\tau$, that is, this greek index labels the flavors of the SM leptons. The $3\times 3$ matrix $V^{\ell}$ is analogous to the Kobayashi-Maskawa quark-mixing matrix KoMa , but for the lepton sector. The set of quantities ${\cal B}_{\alpha\nu_{j}}$ and ${\cal B}_{\alpha N_{j}}$, just defined in Eqs. (42) and (43), respectively define the $3\times 3$ matrices ${\cal B}_{\nu}$ and ${\cal B}_{N}$, which can be accommodated as matrix blocks into the $3\times 6$ matrix
${\cal B}=\big{(}\begin{array}[]{cc}{\cal B}_{\nu}&{\cal B}_{N}\end{array}\big{)}.$ (44)
@@
${\cal B}_{\alpha j}=\left\{\begin{array}[]{lcl}{\cal B}_{\alpha\nu_{k}}\textrm{, if }j=1,2,3,\vspace{0.3cm}\\ {\cal B}_{\alpha N_{k}}\textrm{, if }j=4,5,6,\end{array}\right.$ (45)
@@
${\cal C}=\left(\begin{array}[]{cc}{\cal C}_{\nu\nu}&{\cal C}_{\nu N}\vspace{0.2cm}\\ {\cal C}_{N\nu}&{\cal C}_{NN}\end{array}\right).$ (48)
@@
${\cal C}_{jk}=\left\{\begin{array}[]{ll}{\cal C}_{\nu_{i}\nu_{l}},&\textrm{ if}j=1,2,3\textrm{ and }k=1,2,3,\vspace{0.2cm}\\ {\cal C}_{\nu_{i}N_{l}},&\textrm{ if }j=1,2,3\textrm{ and }k=4,5,6,\vspace{0.2cm}\\ {\cal C}_{N_{i}\nu_{l}},&\textrm{ if }j=4,5,6\textrm{ and }k=1,2,3,\vspace{0.2cm}\\ {\cal C}_{N_{i}N_{l}},&\textrm{ if }j=4,5,6\textrm{ and }k=4,5,6,\end{array}\right.$ (49)
@@
Returning to the Lagrangian ${\cal L}_{\rm BSM}$, as expressed in Eq. (36), the explicit expressions of the Lagrangian terms ${\cal L}_{W\nu l}$, ${\cal L}_{Z\nu\nu}$, and ${\cal L}_{h\nu\nu}$ read
@@
${\cal U}=\left(\begin{array}[]{cc}\big{(}{\bf 1}_{3}+\xi^{*}\xi^{\rm T}\big{)}^{-\frac{1}{2}}&\xi^{*}\big{(}{\bf 1}_{3}+\xi^{\rm T}\xi^{*}\big{)}^{-\frac{1}{2}}\vspace{0.2cm}\\ -\xi^{\rm T}\big{(}{\bf 1}_{3}+\xi^{*}\xi^{\rm T}\big{)}^{-\frac{1}{2}}&\big{(}{\bf 1}_{3}+\xi^{\rm T}\xi^{*}\big{)}^{-\frac{1}{2}}\end{array}\right).$ (58)
@@
${\cal C}=\left(\begin{array}[]{cc}\big{(}{\bf 1}_{3}+\xi\xi^{\dagger}\big{)}^{-1}&\big{(}{\bf 1}_{3}+\xi\xi^{\dagger}\big{)}^{-1}\xi\vspace{0.2cm}\\ \xi^{\dagger}\big{(}{\bf 1}_{3}+\xi\xi^{\dagger}\big{)}^{-1}&\xi^{\dagger}\big{(}{\bf 1}_{3}+\xi\xi^{\dagger}\big{)}^{-1}\xi\end{array}\right).$ (60)